OpenStudy (anonymous):

Which is the side length of a cube with a volume of 2744 cm3?

OpenStudy (anonymous):

I think we got this answer from the last question

OpenStudy (anonymous):

ooooopppppsss

OpenStudy (anonymous):

Which is the side length of a cube with a volume of 2744 cm3?

OpenStudy (anonymous):

Yep, or the cube root of 2744.

OpenStudy (anonymous):

14

sammixboo (sammixboo):

OK so like I said before =) This is how to find the volume of a cube \(\rm s^3~=~V\) V = Volume s = side We know our volume is 2744, so plug it into the formula \(\rm s^3~=~2744\) We know that \(\rm s^3~=~2744\) \(\rightarrow\) \(\rm s~=~2744^{(\dfrac{1}{3})}\) So what is \(\rm 2744^{(\dfrac{1}{3})}\)

OpenStudy (anonymous):